for f(x)=-3x+7, find the value of x that leads to the given value of f(x).
f(x)= 5/2

Answer by stanbon(75887)

for f(x)=-3x+7, find the value of x that leads to the given value of f(x).
f(x)= 5/2
-----
Solve:: -3x+7 = 5/2
-----
-6x + 14 = 5
-6x = -9
x = 3/2
--------------
Cheers,
Stan H.
---------